A rectangle is twice as long as it is wide. If the width is increased by 25% and the length is decreased by 25%, the resulting rectangle will have 200 fewer square units of area than the original rectangle. What is the original width of the rectangle?
Answer by mananth(16946)

A rectangle is twice as long as it is wide. If the width is increased by 25% and the length is decreased by 25%, the resulting rectangle will have 200 fewer square units of area than the original rectangle. What is the original width of the rectangle?
...
width be x
length be 2x
Area = x*2x=2x^2
..
width increase by 25% which means the width becomes 1.25x
length is decreased by 25% which means it becomes 0.75*2x=1.5x
Area = 1.25x*1.5x= 1.875x^2
...
2.0x^2 - 1.875x^2= 200
0.125x^2=200
x^2= 200/0.125
x^2=1600
x= 40 the width
length = 3x = 120 units
...
m.ananth@hotmail.ca